# Existence of distributional solutions to degenerate elliptic systems for locally integrable forcing

### v1

@@ Abstract.
This paper presents an existence result and maximal regularity estimates for distributional solutions to degenerate/singular elliptic systems of $p$-Laplacian type with absorption and (prescribed) locally integrable forcing posed in unbounded Lipschitz domains. In particular, the forcing terms may not belong to the dual space of an energy space, e.g., $W^{1,p}_{\rm loc}$, which is necessary for the existence of weak (or energy) solutions of class $W^{1,p}_{\rm loc}$. The method of a proof relies on both local energy estimates and a relative truncation technique developed by Bulíček and Schwarzacher (Calc. Var. PDEs in 2016), where the bounded domain case is studied for (globally) integrable forcing.
G.A. is supported by JSPS KAKENHI Grant Numbers JP24H00184, JP21KK0044, JP21K18581, JP20H01812 and JP20H00117. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@ 1. Introduction
Let $d,N\in\mathbb{Z}_{\geq 1}$ and let $\Omega$ be an unbounded Lipschitz domain of $\mathbb{R}^{d}$. Let $1<p<r<\infty$ and let $q,s\in(1,\infty)$ satisfy
@@
where $u:\Omega\to\mathbb{R}^{N}$ is an unknown function, and $f\in L^{q}_{\mathrm{loc}}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{s}_{{\mathrm{loc}}}(\Omega;\mathbb{R}^{N})$ are prescribed. In particular, when $\Omega=\mathbb{R}^{d}$, we ignore the Dirichlet boundary condition. Moreover, $A:\Omega\times\mathbb{R}^{d\times N}\to\mathbb{R}^{d\times N}$ is a Carathéodory function complying with the following assumptions: There exist constants $C_{1},C_{2}>0$ and nonnegative functions $\beta_{1}\in L^{1}_{\mathrm{loc}}(\Omega)$, $\beta_{2}\in L^{p^{\prime}}_{\mathrm{loc}}(\Omega)$ such that
@@
In case $f\in L^{p}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{r}(\Omega;\mathbb{R}^{N})$ (i.e., they are globally integrable over $\Omega$ with $p=q$ and $r=s$), existence of weak solutions can be proved in a standard manner (see Proposition 4.1 below). On the other hand, in case either $p>q$ or $r>s$ holds, (P) admits no weak solution for general $f\in L^{q}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{s}(\Omega;\mathbb{R}^{N})$ due to a simple comparison of both sides of the equation. However, we can still expect existence of distributional solutions to (P) (see Definition 2.1 below). The main purpose of the present paper is to prove existence of such distributional solutions to (P) for $f\in L^{q}_{\mathrm{loc}}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{s}_{{\mathrm{loc}}}(\Omega;\mathbb{R}^{N})$ with exponents $q,s\in(1,\infty)$ satisfying (1.1) (whose lower bounds are necessary for the existence of distributional solutions) as well as to establish maximal regularity estimates for such distributional solutions; this is a reason why we consider the forcing term of this form in (P). To be more precise, we expect that $\nabla u$ and $u$ can be estimated in terms of $f$ and $g$ in $L^{q}_{\rm loc}$ and $L^{s}_{\rm loc}$, respectively.
@@
$-\Delta u+|u|^{r-2}u=F\ \mbox{ in }\mathbb{R}^{d}$
@@
for any $R^{\prime}>R>0$. In [3], this result was also extended to degenerate and singular elliptic PDEs of $p$-Laplacian type such as
@@
for $F\in L^{1}_{\rm loc}(\mathbb{R}^{d})$ and $r>p$. On the other hand, the optimal regularity of the (distributional) solution is not fully pursued in these literatures. Actually, the Calderón–Zygmund estimate for the Laplacian fails in $L^{1}$. Moreover, it is rather delicate to determine the optimal regularity of solutions to degenerate and singular nonlinear elliptic PDEs such as (1.5) even posed in bounded domains.
@@
for $f\in L^{q}(\Omega)$ and $q\geq p$. See [11][9, Theorem 4.1][1][13, Theorem 7.8]. Here we also remark that the estimate (1.7) can fail for vectorial cases with $q>p=2$ large enough (see [18]). We further refer the reader to survey papers [15, 16, 17] and references therein. On the other hand, its extension to the case where $1<q<p$ is not straightforward at all in contrast with the linear case, for which a duality argument works. Moreover, even existence of (distributional) solutions is already highly nontrivial when $q<p$. In [12] and [14], maximal regularity estimates are established with $q<p$ close enough to $p$ for distributional solutions defined on $\mathbb{R}^{d}$, provided that they exist. Moreover, an existence result and maximal regularity estimates for distributional solutions to (1.6) are first proved for bounded domains and $q<p$ close enough to $p$ by Bulíček and Schwarzacher in [7], which relies on sophisticated techniques such as an $\mathcal{A}_{p}$-weighted biting div-curl lemma developed in [6, Theorem 2.6] as well as a relative truncation technique for establishing maximal regularity estimates.
The aim of the present paper may also be regarded as an extension of the nonlinear Calderón–Zygmund theory to the elliptic system (P) with an arbitrary locally integrable forcing $f\in L^{q}_{\rm loc}(\Omega;\mathbb{R}^{d\times N})$ for $q$ being in the range $1<q<p$. Here we emphasize again that the presence of the absorption term in (P) is necessary to prove an existence result as mentioned above (see [4, 3]), and furthermore, the forcing terms of the equation in (P) are designed in such a way as to discuss the maximal regularity of gradients of distributional solutions.
@@ 2. Main Result
for any $R>0$. In what follows, the homogeneous Dirichlet boundary condition for $u\in W^{1,1}_{\rm loc}(\Omega;\mathbb{R}^{N})$ will be regarded (in a weak sense) as
@@
Let $1<p<r<\infty$ and let $f\in L^{\max\{1,p-1\}}_{\rm loc}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{\max\{1,r-1\}}_{\rm loc}(\Omega;\mathbb{R}^{N})$.
@@
    A function $u\in W_{\mathrm{loc}}^{1,\max\{1,p-1\}}(\Omega;\mathbb{R}^{N})\cap L_{\mathrm{loc}}^{\max\{1,r-1\}}(\Omega;\mathbb{R}^{N})$ is called a distributional solution to (P), if it holds that
@@
Let $\Omega$ be an unbounded Lipschitz domain of $\mathbb{R}^{d}$ and let $1<p<r<\infty$. Then there exists a positive constant $\varepsilon_{0}=\varepsilon_{0}(C_{1},C_{2},\Omega,d,N,p,r)\in(0,p/r]$ satisfying the following: Let $\varepsilon\in(0,\varepsilon_{0})$ be fixed and set
@@
Then for any $f\in L_{\mathrm{loc}}^{q}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L_{\mathrm{loc}}^{s}(\Omega;\mathbb{R}^{N})$, the Dirichlet problem (P) admits a distributional solution $u\in W^{1,q}_{\mathrm{loc}}(\Omega;\mathbb{R}^{N})\cap L^{s}_{\mathrm{loc}}(\Omega;\mathbb{R}^{N})$ such that
@@ 3.1. Muckenhoupt weights
We next set up a couple of lemmas for later use. The following lemma is concerned with some relations between maximal functions and Muckenhoupt weights. We refer the reader to [19, pp. 229–230] and [20, p. 5] (see also [7, Lemma 2.1]) for more details.
@@
Let us recall the following lemma, which is concerned with continuous embeddings from $\mathcal{A}_{p}$-weighted Lebesgue spaces to usual ones (see [6, pp. 1125,1126]).
@@ 3.2. $\mathcal{A}_{p}$-weighted biting div-curl lemma
In this subsection, we recall the $\mathcal{A}_{p}$-weighted biting div-curl lemma developed in [6, Theorem 2.6].
Lemma 3.5 ($\mathcal{A}_{p}$-weighted biting div-curl lemma [6, Theorem 2.6]).
@@
The assumption (ii) is stronger than the original one of [6, Theorem 2.6], but it is enough to apply the lemma in the present paper.
@@ 3.3. Whitney decomposition of open sets
We refer the reader to [8, Proposition 3.17] and [7, Proposition 3.1] for proofs of (i)–(iv) and (v)–(vi), respectively.
@@ 5.2. Weighted local energy estimates
Developing the argument used in [7, Proposition 4.1], we shall prove the following proposition, which will play a key role for proving Theorem 2.2.
@@
Let $\Omega$ be an unbounded Lipschitz domain of $\mathbb{R}^{d}$ and let $1<p<r<\infty$. Then there exists a constant $\varepsilon_{1}=\varepsilon_{1}(C_{1},C_{2},\Omega,d,N,p,r)>0$ satisfying the following: Let $f\in L^{p}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{r}(\Omega;\mathbb{R}^{N})$ and let $u\in X^{p,r}(\Omega;\mathbb{R}^{N})$ be the unique weak solution to (P). Let $\varepsilon\in(0,\varepsilon_{1})$ and $h\in L^{1}_{{\mathrm{loc}}}(\Omega)\setminus\{0\}$ be arbitrarily fixed. Then there exists a constant $C_{0}=C_{0}(C_{1},C_{2},\Omega,d,N,p,r,\varepsilon)>0$ such that
@@
where $\omega$ is the weight function given by $\omega=(M[\bar{h}\chi_{\Omega_{2R}}]+\delta)^{-\varepsilon}$ with the zero extension $\bar{h}$ of $h$ onto $\mathbb{R}^{d}$ and $\beta_{1}\in L^{1}_{\mathrm{loc}}(\Omega)$ and $\beta_{2}\in L^{p^{\prime}}_{\mathrm{loc}}(\Omega)$ are nonnegative functions satisfying (1.2) and (1.3), for every $R>0$ and $\delta\in(0,1]$.
@@
We can estimate $J_{1}$ as in the proof of [7, Proposition 4.1], but we also give a detail for the completeness. Set $G:=|f|^{p-1}+C_{2}|\nabla u|^{p-1}+\beta_{2}$. For each $\alpha\in(0,1)$ fixed, using Proposition 3.6 and noting that $\sum_{j\in A_{i}}\psi_{j}\equiv 1$ on $Q_{i}$, one can verify that
@@
Here we employed [7, Lemma 3.1] along with (vi) of Proposition 3.6, in particular, $\mathrm{diam}\,(Q_{i})|\nabla\psi_{i}|\leq c(d)$, in order to derive the first inequality ()(**)( ∗ ∗ ). Moreover, we also used the fact that $(3/2)Q_{j}\subset 6Q_{i}$ if $j\in A_{i}$. Now, from (ii) of Proposition 3.6, we find that $9Q_{i}\cap\mathcal{O}_{\lambda}^{c}\neq\emptyset$, and hence, for a.e. $x_{0}\in 9Q_{i}\cap\mathcal{O}_{\lambda}^{c}$, it follows from the definition of $\mathcal{O}_{\lambda}$ that
@@
where $\varepsilon_{1}$ is the constant given in Proposition 5.3. Let $\varepsilon\in(0,\varepsilon_{0})$ be fixed arbitrarily. Set $q:=p-\varepsilon\in(p-1,p)$ and $s:=r(p-\varepsilon)/p\in(r-1,r)$. Moreover, let $f\in L_{{\mathrm{loc}}}^{q}(\Omega;\mathbb{R}^{d\times N})$, $g\in L_{{\mathrm{loc}}}^{s}(\Omega;\mathbb{R}^{N})$ and fix $R>0$.
@@
$\delta:=1\quad\mbox{ and }\quad h:=|f|+|g|^{s/q}\in L^{q}_{\rm loc}(\Omega).$
@@
Thanks to Lemma 3.4, since $\omega^{-1}=(M[h\chi_{\Omega_{2R}}]+1)^{p-q}\in L^{q/(p-q)}_{\rm loc}(\mathbb{R}^{d})=L^{s/(r-s)}_{\rm loc}(\mathbb{R}^{d})$ (indeed, $h\in L^{q}_{\rm loc}(\Omega)$ and $M$ is bounded in $L^{q}(\mathbb{R}^{d})$) and $q/(p-q)=s/(r-s)$ from $s=rq/p$, both embeddings $L^{p}_{\omega}(\Omega_{R};\mathbb{R}^{d\times N})\hookrightarrow L^{q}(\Omega_{R};\mathbb{R}^{d\times N})$ and $L^{r}_{\omega}(\Omega_{R};\mathbb{R}^{N})\hookrightarrow L^{s}(\Omega_{R};\mathbb{R}^{N})$ are continuous. Therefore $(\nabla u^{k})_{k}$ is bounded in $L^{q}(\Omega_{R};\mathbb{R}^{d\times N})$, and $(u^{k})_{k}$ is so in $L^{s}(\Omega_{R};\mathbb{R}^{N})$. {comment} Thus it only remains to show the boundedness of $(u^{k})_{k}$ in $L^{s}(\Omega_{R};\mathbb{R}^{N})$ as $k\to\infty$. Let $\varphi\in L^{s^{\prime}}(\Omega_{R};\mathbb{R}^{N})$ and note that $q/(p-q)\geq s/(r-s)$ from $s\leq rq/p$. Since $\omega^{-1}\in L^{q/(p-q)}(\Omega_{R})\subset L^{s/(r-s)}(\Omega_{R})$, one has $\omega^{1-r^{\prime}}\in L^{s/((r-s)(r^{\prime}-1))}(\Omega_{R})$. Hence
@@
where $1<q<p<r<\infty$, $1<s<r$, $f\in L^{q}_{\mathrm{loc}}(\Omega;\mathbb{R}^{d\times N})$, $g\in L^{s}_{\mathrm{loc}}(\Omega;\mathbb{R}^{N})$ and $A:\Omega\times\mathbb{R}^{d\times N}\to\mathbb{R}^{d\times N}$ is given as before. Moreover, $\beta:\Omega\times\mathbb{R}^{N}\to\mathbb{R}^{N}$ is also supposed to be a Carathéodory function satisfying the following assumptions: There exist constants $C_{3},C_{4}>0$ and nonnegative functions $\beta_{3}\in L^{1}_{\mathrm{loc}}(\Omega)$, $\beta_{4}\in L^{r^{\prime}}_{\mathrm{loc}}(\Omega)$ such that

### v2

@@ Abstract.
This paper presents an existence result and maximal regularity estimates for distributional solutions to degenerate/singular elliptic systems of $p$-Laplacian type with absorption and (prescribed) locally integrable forcing posed in (possibly unbounded) Lipschitz domains. In particular, the forcing terms may not belong to the dual space of an energy space, e.g., $W^{1,p}_{\rm loc}$, which is necessary for the existence of weak (or energy) solutions of class $W^{1,p}_{\rm loc}$. The method of a proof relies on both local energy estimates and a relative truncation technique developed by Bulíček and Schwarzacher (Calc. Var. PDEs in 2016), where the bounded domain case is studied for (globally) integrable forcing.
@@ 1. Introduction
Let $d,N\in\mathbb{Z}_{\geq 1}$ and let $\Omega$ be a (possibly unbounded) Lipschitz domain of $\mathbb{R}^{d}$. Let $1<p<r<\infty$ and let $q,s\in(1,\infty)$ satisfy
@@
where $u:\Omega\to\mathbb{R}^{N}$ is an unknown function, and $f\in L^{q}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{d\times N})$ and $g\in L^{s}_{{\mathrm{loc}}}(\overline{\Omega};\mathbb{R}^{N})$ are prescribed. In particular, when $\Omega=\mathbb{R}^{d}$, we ignore the Dirichlet boundary condition. Moreover, $A:\Omega\times\mathbb{R}^{d\times N}\to\mathbb{R}^{d\times N}$ is a Carathéodory function complying with the following assumptions: There exist constants $C_{1},C_{2}>0$ and nonnegative functions $\beta_{1}\in L^{1}_{\mathrm{loc}}(\overline{\Omega})$, $\beta_{2}\in L^{p^{\prime}}_{\mathrm{loc}}(\overline{\Omega})$ such that
@@
In case $f\in L^{p}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{r}(\Omega;\mathbb{R}^{N})$ (i.e., they are globally integrable over $\Omega$ with $p=q$ and $r=s$), existence of weak solutions can be proved in a standard manner (see Proposition 4.1 below). On the other hand, in case either $p>q$ or $r>s$ holds, (P) admits no weak solution for general $f\in L^{q}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{s}(\Omega;\mathbb{R}^{N})$ due to a simple comparison of both sides of the equation. However, we can still expect existence of distributional solutions to (P) (see Definition 2.1 below). The main purpose of the present paper is to prove existence of such distributional solutions to (P) for $f\in L^{q}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{d\times N})$ and $g\in L^{s}_{{\mathrm{loc}}}(\overline{\Omega};\mathbb{R}^{N})$ with exponents $q,s\in(1,\infty)$ satisfying (1.1) (whose lower bounds are necessary for the existence of distributional solutions) as well as to establish maximal regularity estimates for such distributional solutions; this is a reason why we consider the forcing term of this form in (P). To be more precise, we expect that $\nabla u$ and $u$ can be estimated in terms of $f$ and $g$ in $L^{q}_{\rm loc}$ and $L^{s}_{\rm loc}$, respectively.
@@
$-\Delta u+|u|^{r-2}u=F\ \mbox{ in }\mathbb{R}^{d}$ (1.5)
@@
for any $R^{\prime}>R>0$. Such an elliptic equation (1.5) with an absorption naturally appears as a stationary problem for the heat equation with an absorption term. Moreover, [4] may also be inspired by studies on existence of solutions to parabolic equations with growing initial data (see, e.g., [12]). Actually, in [4], a parabolic counterpart of the result for (1.5) is also briefly discussed (e.g., solvability of the heat equation with an absorption is considered for any locally integrable initial data). In [3], the aforementioned result is also extended to degenerate and singular elliptic PDEs of $p$-Laplacian type such as
@@
for $F\in L^{1}_{\rm loc}(\mathbb{R}^{d})$ and $r>p$. The above case is much more technically involved than the semilinear case. Indeed, in contrast with [4], one has to establish a gradient estimate for certain approximate solutions $(u_{n})$ due to the nonlinearity arising from the $p$-Laplacian. It is actually done in [3, Lemma 2.2] but in a totally different way from the present paper. Moreover, in order to identify a weak limit of the $p$-Laplacian term $\Delta_{p}u_{n}$, a pointwise convergence of the gradients $\nabla u_{n}$ is also derived by use of a strong monotonicity of the $p$-Laplacian (see assumption (A4) of [3]), which is no longer assumed in the present paper (cf. see (1.4)).
On the other hand, the optimal regularity of the (distributional) solution is not fully pursued in these literatures. Actually, the Calderón–Zygmund estimate for the Laplacian fails in $L^{1}$. Moreover, it is rather delicate to determine the optimal regularity of solutions to degenerate and singular nonlinear elliptic PDEs such as (1.6) even posed in bounded domains.
@@
for $f\in L^{q}(\Omega)$ and $q\geq p$. See [15][13, Theorem 4.1][1][17, Theorem 7.8]. Here we also remark that the estimate (1.8) can fail for vectorial cases with $q>p=2$ large enough (see [23]). We further refer the reader to survey papers [20, 21, 22] and references therein. On the other hand, its extension to the case where $1<q<p$ is not straightforward at all in contrast with the linear case, for which a duality argument works. Moreover, even existence of (distributional) solutions is already highly nontrivial when $q<p$. In [16] and [18], maximal regularity estimates are established with $q<p$ close enough to $p$ for distributional solutions defined on $\mathbb{R}^{d}$, provided that they exist. Moreover, an existence result and maximal regularity estimates for distributional solutions to (1.7) are first proved for bounded domains and $q<p$ close enough to $p$ by Bulíček and Schwarzacher in [10], which relies on sophisticated techniques such as an $\mathcal{A}_{p}$-weighted biting div-curl lemma developed in [9, Theorem 2.6] as well as a relative truncation technique for establishing maximal regularity estimates.
The aim of the present paper may also be regarded as an extension of the nonlinear Calderón–Zygmund theory to the elliptic system (P) with an arbitrary locally integrable forcing $f\in L^{q}_{\rm loc}(\overline{\Omega};\mathbb{R}^{d\times N})$ and $g\in L^{s}_{\rm loc}(\overline{\Omega};\mathbb{R}^{N})$ for $q$ and $s$ being in the range (1.1). Here we emphasize again that the presence of the absorption term in (P) is necessary to prove an existence result as mentioned above (see [4, 3]), and furthermore, the forcing terms of the equation in (P) are designed in such a way as to discuss the maximal regularity of gradients of distributional solutions.
Finally, we give a couple of related references without any claim of the completeness, although all these results are devoted to equations of different type posed in bounded domains only. The results in [9, 10] have been extended to stationary Navier–Stokes equations with non-Newtonian stress tensors in [19], where a solenoidal Lipschitz truncation is also developed (cf. see also [6], which is a generalization of [9]). One can find parabolic counterparts of these studies in, e.g., [7, 8] and references therein.
@@ 2. Main Result
for any $R>0$. In what follows, the homogeneous Dirichlet boundary condition for $u\in W^{1,1}_{\rm loc}(\overline{\Omega};\mathbb{R}^{N})$ will be regarded (in a weak sense) as
@@
Let $1<p<r<\infty$ and let $f\in L^{\max\{1,p-1\}}_{\rm loc}(\overline{\Omega};\mathbb{R}^{d\times N})$ and $g\in L^{\max\{1,r-1\}}_{\rm loc}(\overline{\Omega};\mathbb{R}^{N})$.
@@
    A function $u\in W_{\mathrm{loc}}^{1,\max\{1,p-1\}}(\overline{\Omega};\mathbb{R}^{N})\cap L_{\mathrm{loc}}^{\max\{1,r-1\}}(\overline{\Omega};\mathbb{R}^{N})$ is called a distributional solution to (P), if it holds that
@@
Let $\Omega$ be a (possibly unbounded) Lipschitz domain of $\mathbb{R}^{d}$ and let $1<p<r<\infty$. Then there exists a positive constant $\varepsilon_{0}=\varepsilon_{0}(C_{1},C_{2},\Omega,d,N,p,r)\in(0,p/r]$ satisfying the following: Let $\varepsilon\in(0,\varepsilon_{0})$ be fixed and set
@@
Then for any $f\in L_{\mathrm{loc}}^{q}(\overline{\Omega};\mathbb{R}^{d\times N})$ and $g\in L_{\mathrm{loc}}^{s}(\overline{\Omega};\mathbb{R}^{N})$, the Dirichlet problem (P) admits a distributional solution $u\in W^{1,q}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{N})\cap L^{s}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{N})$ such that
@@ 3.1. Muckenhoupt weights
We next set up a couple of lemmas for later use. The following lemma is concerned with some relations between maximal functions and Muckenhoupt weights. We refer the reader to [24, pp. 229–230] and [25, p. 5] (see also [10, Lemma 2.1]) for more details.
@@
Let us recall the following lemma, which is concerned with continuous embeddings from $\mathcal{A}_{p}$-weighted Lebesgue spaces to usual ones (see [9, pp. 1125,1126]).
@@ 3.2. $\mathcal{A}_{p}$-weighted biting div-curl lemma
In this subsection, we recall the $\mathcal{A}_{p}$-weighted biting div-curl lemma developed in [9, Theorem 2.6].
Lemma 3.5 ($\mathcal{A}_{p}$-weighted biting div-curl lemma [9, Theorem 2.6]).
@@
The assumption (ii) is stronger than the original one of [9, Theorem 2.6], but it is enough to apply the lemma in the present paper.
@@ 3.3. Whitney decomposition of open sets
We refer the reader to [11, Proposition 3.17] and [10, Proposition 3.1] for proofs of (i)–(iv) and (v)–(vi), respectively.
@@ 5.2. Weighted local energy estimates
Developing the argument used in [10, Proposition 4.1], we shall prove the following proposition, which will play a key role for proving Theorem 2.2.
@@
Let $\Omega$ be a (possibly unbounded) Lipschitz domain of $\mathbb{R}^{d}$ and let $1<p<r<\infty$. Then there exists a constant $\varepsilon_{1}=\varepsilon_{1}(C_{1},C_{2},\Omega,d,N,p,r)>0$ satisfying the following: Let $f\in L^{p}(\Omega;\mathbb{R}^{d\times N})$ and $g\in L^{r}(\Omega;\mathbb{R}^{N})$ and let $u\in X^{p,r}(\Omega;\mathbb{R}^{N})$ be the unique weak solution to (P). Let $\varepsilon\in(0,\varepsilon_{1})$ and $h\in L^{1}_{{\mathrm{loc}}}(\overline{\Omega})\setminus\{0\}$ be arbitrarily fixed. Then there exists a constant $C_{0}=C_{0}(C_{1},C_{2},\Omega,d,N,p,r,\varepsilon)>0$ such that
@@
where $\omega$ is the weight function given by $\omega=(M[\bar{h}\chi_{\Omega_{2R}}]+\delta)^{-\varepsilon}$ with the zero extension $\bar{h}$ of $h$ onto $\mathbb{R}^{d}$ and $\beta_{1}\in L^{1}_{\mathrm{loc}}(\overline{\Omega})$ and $\beta_{2}\in L^{p^{\prime}}_{\mathrm{loc}}(\overline{\Omega})$ are nonnegative functions satisfying (1.2) and (1.3), for every $R>0$ and $\delta\in(0,1]$.
@@
We can estimate $J_{1}$ as in the proof of [10, Proposition 4.1], but we also give a detail for the completeness. Set $G:=|f|^{p-1}+C_{2}|\nabla u|^{p-1}+\beta_{2}$. For each $\alpha\in(0,1)$ fixed, using Proposition 3.6 and noting that $\sum_{j\in A_{i}}\psi_{j}\equiv 1$ on $Q_{i}$, one can verify that
@@
Here we employed [10, Lemma 3.1] along with (vi) of Proposition 3.6, in particular, $\mathrm{diam}\,(Q_{i})|\nabla\psi_{i}|\leq c(d)$, in order to derive the first inequality ()(**)( ∗ ∗ ). Moreover, we also used the fact that $(3/2)Q_{j}\subset 6Q_{i}$ if $j\in A_{i}$. Now, from (ii) of Proposition 3.6, we find that $9Q_{i}\cap\mathcal{O}_{\lambda}^{c}\neq\emptyset$, and hence, for a.e. $x_{0}\in 9Q_{i}\cap\mathcal{O}_{\lambda}^{c}$, it follows from the definition of $\mathcal{O}_{\lambda}$ that
@@
where $\varepsilon_{1}$ is the constant given in Proposition 5.3. Let $\varepsilon\in(0,\varepsilon_{0})$ be fixed arbitrarily. Set $q:=p-\varepsilon\in(p-1,p)$ and $s:=r(p-\varepsilon)/p\in(r-1,r)$. Moreover, let $f\in L_{{\mathrm{loc}}}^{q}(\overline{\Omega};\mathbb{R}^{d\times N})$, $g\in L_{{\mathrm{loc}}}^{s}(\overline{\Omega};\mathbb{R}^{N})$ and fix $R>0$.
@@
$\delta:=1\quad\mbox{ and }\quad h:=|f|+|g|^{s/q}\in L^{q}_{\rm loc}(\overline{\Omega}).$
@@
Thanks to Lemma 3.4, since $\omega^{-1}=(M[h\chi_{\Omega_{2R}}]+1)^{p-q}\in L^{q/(p-q)}_{\rm loc}(\mathbb{R}^{d})=L^{s/(r-s)}_{\rm loc}(\mathbb{R}^{d})$ (indeed, $h\in L^{q}_{\rm loc}(\overline{\Omega})$ and $M$ is bounded in $L^{q}(\mathbb{R}^{d})$) and $q/(p-q)=s/(r-s)$ from $s=rq/p$, both embeddings $L^{p}_{\omega}(\Omega_{R};\mathbb{R}^{d\times N})\hookrightarrow L^{q}(\Omega_{R};\mathbb{R}^{d\times N})$ and $L^{r}_{\omega}(\Omega_{R};\mathbb{R}^{N})\hookrightarrow L^{s}(\Omega_{R};\mathbb{R}^{N})$ are continuous. Therefore $(\nabla u^{k})_{k}$ is bounded in $L^{q}(\Omega_{R};\mathbb{R}^{d\times N})$, and $(u^{k})_{k}$ is so in $L^{s}(\Omega_{R};\mathbb{R}^{N})$. {comment} Thus it only remains to show the boundedness of $(u^{k})_{k}$ in $L^{s}(\Omega_{R};\mathbb{R}^{N})$ as $k\to\infty$. Let $\varphi\in L^{s^{\prime}}(\Omega_{R};\mathbb{R}^{N})$ and note that $q/(p-q)\geq s/(r-s)$ from $s\leq rq/p$. Since $\omega^{-1}\in L^{q/(p-q)}(\Omega_{R})\subset L^{s/(r-s)}(\Omega_{R})$, one has $\omega^{1-r^{\prime}}\in L^{s/((r-s)(r^{\prime}-1))}(\Omega_{R})$. Hence
@@
where $1<q<p<r<\infty$, $1<s<r$, $f\in L^{q}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{d\times N})$, $g\in L^{s}_{\mathrm{loc}}(\overline{\Omega};\mathbb{R}^{N})$ and $A:\Omega\times\mathbb{R}^{d\times N}\to\mathbb{R}^{d\times N}$ is given as before. Moreover, $\beta:\Omega\times\mathbb{R}^{N}\to\mathbb{R}^{N}$ is also supposed to be a Carathéodory function satisfying the following assumptions: There exist constants $C_{3},C_{4}>0$ and nonnegative functions $\beta_{3}\in L^{1}_{\mathrm{loc}}(\overline{\Omega})$, $\beta_{4}\in L^{r^{\prime}}_{\mathrm{loc}}(\overline{\Omega})$ such that
@@
Acknowledgments
The authors wish to thank anonymous referees for their valuable comments. G.A. is supported by JSPS KAKENHI Grant Numbers JP24H00184, JP21KK0044, JP21K18581, JP20H01812 and JP20H00117. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.